QS= a+bp, QD = c-hp, what is equilibrium price and quantity demanded ?


1
Expert's answer
2021-11-28T17:48:24-0500

At equilibrium

Quantity supplied equal quantity demanded.


"Q_s =Q_d"


"\therefore"


"a+bp= c-hp"


"hp+bp = c-a"


"p(h+b)= c-a"


"P_e" equilibrium price is


"p_e= \frac{c-a}{h+b}"


"Q_s = Q_d"


Quantity at equilibrium"Q_e"


"Q_e = \frac{Q_s}{Q_d}"


"Q_e = \frac{a+bp}{c-hp}"